# Human Arm Pose Estimation with a Shoulder-worn Force-Myography Device for Human-Robot Interaction

### v1

@@ I Introduction
In this letter, we explore the feasibility of utilizing FMG for comprehensive modeling of the human arm, with the goal of natural HRI. In a novel approach, FMG sensors are strategically placed on the upper back, shoulder and arm, enabling the acquisition of data that can be used to infer arm pose. This is the first application of FMG on shoulder muscles and for modeling the state of the entire human arm. To estimate the instantaneous pose, we propose a novel approach utilizing the Transformer architecture [32], leveraging its ability to process temporal sequential data effectively. We then demonstrate its use in an HRI scenario where a robot must perform its own task without interfering or colliding with the human user. Unlike the relative positioning of IMUs, FMG offers absolute positioning with respect to the human torso without the need for constant calibration. Hence, it can be integrated in the future to the clothing of users, enabling seamless sensing and data collection in various applications.
@@ II-C Data Collection
Train and test datasets were collected by synchronously recording FMG signals with the three key points on the human arm. This research investigates the feasibility of training an accurate $\Gamma$ model using data from only a single participant and potentially transferring it to novel users. To achieve this, data was collected from a single participant across $K$ separate sessions and $M$ samples per session. In each session, the device was taken off and re-worn to include variations in sensor placement and tightening forces.
@@ II-D Pose Estimation Model
Our model, illustrated in Figure 3 consists of an embedding layer that maps the input data into a space of dimension $d_{model}$. A positional encoding layer is applied to inject information about the sequence order into the model. The model includes a transformer encoder with multiple layers, each comprising a multi-head attention mechanism, a feed-forward network with $d_{ff}$ hidden layers and ReLU activation after each layer. Dropout is applied after each attention and feed-forward sub-layer. The output of the Transformer is passed through a series of fully connected layers, each reducing the dimensionality of the feature space, yielding the final position estimation outputs $\tilde{\mathbf{p}}_{el}$ and $\tilde{\mathbf{p}}_{wr}$. In addition, the output of the Transformer is passed through another fully-connected network outputting $\mathbf{x}^{\prime}$, the reconstruction of the FMG input $\mathbf{x}$. The Transformer-based model is trained in two phases [32]. In the first phase, the model is trained in an unsupervised manner where the positional outputs are ignored and the reconstruction loss $\mathcal{L}=\|\mathbf{x}-\mathbf{x}^{\prime}\|$ is minimized. In the second phase, supervised learning is employed to minimize the loss given by
@@ III-A Dataset
Figure 4: Real and predicted positions of the (top) elbow and (bottom) wrist, with regard to motion time.
@@ III-B Model Evaluation
In this section, we analyse the Transformer-based model. First, we compare our proposed model to other existing models including Fully-Connected Neural Network (FC-NN), 1DCNN [36], DLinear [37] and Long Short-Term Memory with CNN (CNN-LSTM) [38].
The FC-NN is the baseline where the temporal data is flattened to a $n\times H$ vector and fed into a series of fully-connected layers. 1DCNN is a type of CNN specifically designed for processing sequential data, such as text or time series. It applies filters to input sequences to extract relevant features and learn patterns. DLinear is a deep linear model designed for forecasting long-term time series and combines the strengths of linear models with the flexibility of deep architectures. It uses a linear transformation to capture temporal dependencies while maintaining computational efficiency. Lastly, CNN-LSTM combines CNNs for feature extraction and LSTM for sequence processing. It passes the data through a series of 1D convolutional layers followed by a sequence of LSTM layers and an FC layer.
@@
Table I summarizes the comparative analysis between the models. The results show the Root Mean Square Error (RMSE) for the elbow and wrist positions over the test data and the mean inference time for a single prediction. For both elbow and wrist, the Transformer provides predictions with the lowest error. Predicting wrist position generally poses more challenges than elbow position due to the greater complexity of wrist movements. This is reflected in the slightly larger error values observed for wrist position estimation. While the Transformer architecture offers better performance compared to other models, it generally requires slightly longer inference times due to its increased complexity. However, the inference time remains sufficiently low and enables real-time applications. Figure 4 shows an example of the predicted elbow and wrist positions along some arm trajectory, based on FMG measurements. The mean position errors for the elbow and wrist along the example path are 65.4 mm and 116.6 mm, respectively.
@@ III-C Feature Importance
where $e$ is the mean error of the non-permuted model and $e_{i}$ is the mean error when feature $i$ is permuted. Table II presents the feature importance scores for each FSR sensor on the user’s arm, averaged across wrist and elbow position errors. Sensor placements and importance score heatmap are illustrated in Figure 6. Out of the 32 sensors, 11 were found to be particularly influential, with importance scores exceeding 3%. This suggests that entire rows of FSR sensors can be removed and a more compact FMG device, utilizing a reduced number of sensors, could be designed without significant compromises in accuracy.
TABLE II: Feature importance scores of the FSR sensors on the FMG device.
@@ III-D Model Transfer
To evaluate the model’s generalization capabilities, we tested it on two new participants with different Chest Circumferences (CC), Shoulder Widths (SW) and Upper-Arm Circumferences (UAC), who were not included in the original training dataset. Data was collected from these participants following the same procedure outlined in Section II-C, each having 10,000 labeled samples yielding 4,850 temporal sequences. Table III summarizes the results along with anthropometric measures of all participants including the one used for training. While the model’s accuracy decreased slightly when tested on new participants with different anthropometric measurements, the results demonstrate its ability to generalize and perform effectively. To improve transferability, one can collect data from various users or fine-tune the above model with limited data from the new user.
TABLE III: Accuracy of model transfer to new users
@@
Train 98 44 28 61 122
1 99 46 30.5 100 190
2 92 44 25.5 95 186
TABLE IV: HRI demonstration success rates
@@ III-E Real-time HRI Demonstration
Two experimental sessions were conducted. In the first session, the user deliberately interfered the motion of the robot by blocking its path. An example of such trial is demonstrated in Figure 7. In the second session, the user naturally performed various tasks, such as picking up and placing tools and other objects scattered on the table (Figures 8-9), while blocking the path of the robot. Each of the sessions consisted of 60 trials. Table IV presents the success rate for the two sessions, including metrics for false positive with failure to stop and false negatives with unnecessary stops. For the latter, a momentary large error made the robot estimate a potential collision leading to a stop. Nevertheless, the majority of trials resulted in true positives, demonstrating the potential of this approach for natural HRI, without visual perception.
@@ IV Conclusions
We investigated the use of a low-cost wearable FMG device for human arm pose estimation in Human-Robot Interaction (HRI) scenarios. This device, worn on the user’s shoulder, incorporates 32 FSR sensors. Using a Transformer-based model, FMG measurements are mapped to the positions of the elbow and wrist, with respect to the shoulder. Our approach exhibited the ability to provide robust and fairly accurate arm pose estimations. Unlike common IMU solutions for arm pose estimation, the FMG device can provide drift-free and absolute estimations without the need for frequent calibrations. The trained model was also shown to be transferable to new users with slight decline in accuracy. Furthermore, a set of HRI experiments has demonstrated the robot’s ability to avoid collisions using the estimated arm poses in a shared workspace. Generally, the FMG device has the potential to enhance various HRI applications, including medical procedures, factory work and assistive technologies.
While the estimation model was shown to be able to fairly generalize to new users, future work should consider making the model more robust to new users. One approach to enhance the model’s generalization capabilities is to train it on a diverse dataset encompassing data from multiple users, as demonstrated in [31]. Alternatively, future work could involve customizing both the FMG device and the model to individual users. This could include 3D scanning a user’s body to create a custom-fitted device, ensuring optimal sensor placement and data quality. Then, the model can be customized for the user through fine-tuning with a limited amount of additional data, as in [26]. In further work, the integration of FSR sensors into wearable textiles could open up new possibilities for advanced and more spontaneous HRI. Future work could also explore the integration of FMG with visual perception to create a robust hybrid system capable of handling diverse environmental conditions.

### v2

@@ I Introduction
In this letter, we explore the feasibility of utilizing FMG for comprehensive modeling of the human arm, with the goal of natural HRI. In a novel approach, FMG sensors are strategically placed on the upper back, shoulder and arm, enabling the acquisition of data that can be used to infer arm pose. This is the first application of FMG on shoulder muscles and for modeling the state of the entire human arm. To estimate the instantaneous pose, we propose a novel approach utilizing the Transformer architecture [32], leveraging its ability to process temporal sequential data effectively. We then demonstrate its use in an HRI scenario where a robot must perform its own task without interfering or colliding with the human user. Unlike the relative positioning of IMUs, FMG offers absolute positioning with respect to the human torso without the need for constant calibration. Also, in contrast to visual perception, it is environment-agnostic and does not rely on line-of-sight or ambient lighting conditions. Hence, it can be integrated in the future to the clothing of users, enabling seamless sensing and data collection in various applications.
@@ II-C Data Collection
Train and test datasets were collected by synchronously recording FMG signals with the three key points on the human arm. This research investigates the feasibility and requirements of training a $\Gamma$ model using data from only a single participant and potentially transferring it to novel users. To achieve this, data was collected from a one participant across $K$ separate sessions and $M$ samples per session. In each session, the device was taken off and re-worn to include variations in sensor placement and tightening forces.
@@ II-D Pose Estimation Model
Our model, illustrated in Figure 3 consists of an embedding layer that maps the input data into a space of dimension $d_{model}$. A positional encoding layer is applied to inject information about the sequence order into the model. The model includes a transformer encoder with multiple layers, each comprising a multi-head attention mechanism, a feed-forward network with $d_{ff}$ hidden layers and ReLU activation after each layer. Dropout is applied after each attention and feed-forward sub-layer. The output of the Transformer is passed through a series of Fully Connected (FC) layers, each reducing the dimensionality of the feature space, yielding the final position estimation outputs $\tilde{\mathbf{p}}_{el}$ and $\tilde{\mathbf{p}}_{wr}$. In addition, the output of the Transformer is passed through another fully-connected network outputting $\mathbf{x}^{\prime}$, the reconstruction of the FMG input $\mathbf{x}$. The Transformer-based model is trained in two phases [32]. In the first phase, the model is trained in an unsupervised manner where the positional outputs are ignored and the reconstruction loss $\mathcal{L}=\|\mathbf{x}-\mathbf{x}^{\prime}\|$ is minimized. In the second phase, supervised learning is employed to minimize the loss given by
@@ III-A Dataset
Figure 4: Real and predicted positions of the (top) elbow and (bottom) wrist, with regard to motion time. The mean position errors for the elbow and wrist along the example path are 65.4 mm and 116.6 mm, respectively.
@@ III-B Model Evaluation
In this section, we analyse the Transformer-based model. First, we compare our proposed model to other existing models including Fully-Connected Neural Network (FC-NN), 1DCNN [36], DLinear [37] and Long Short-Term Memory with CNN (CNN-LSTM) [38]. The FC-NN is the baseline where the temporal data is flattened to a $n\times H$ vector and fed into a series of fully-connected layers. 1DCNN is a type of CNN specifically designed for processing sequential data, such as text or time series. It applies filters to input sequences to extract relevant features and learn patterns. DLinear is a deep linear model designed for forecasting long-term time series and combines the strengths of linear models with the flexibility of deep architectures. It uses a linear transformation to capture temporal dependencies while maintaining computational efficiency. Lastly, CNN-LSTM combines CNNs for feature extraction and LSTM for sequence processing. It passes the data through a series of 1D convolutional layers followed by a sequence of LSTM layers and an FC layer.
@@
Table I summarizes the comparative analysis between the models. The results show the Root Mean Square Error (RMSE) for the elbow and wrist positions over the test data and the mean inference time for a single prediction. For both elbow and wrist, the Transformer provides predictions with the lowest error. While the elbow estimation is influenced only by the motions of the shoulder, the wrist’s position is more affected by both the elbow and shoulder angles and, thus, is more susceptible to errors propagated through the kinematic chain. This is reflected in the larger error values observed for wrist position estimation compared to the elbow estimations. While the Transformer architecture offers better performance compared to other models, it generally requires slightly longer inference times due to its increased complexity. However, the inference time remains sufficiently low and enables real-time applications. Figure 4 shows an example of the predicted elbow and wrist positions along some arm trajectory, based on FMG measurements.
@@ III-C Feature Importance
where $e$ is the mean error of the non-permuted model and $e_{i}$ is the mean error when feature $i$ is permuted. Table II presents the feature importance scores for each FSR sensor on the user’s arm, averaged across wrist and elbow position errors. Sensor placements and importance score heatmap are illustrated in Figure 6. Out of the 32 sensors, 11 were found to be particularly influential, with importance scores exceeding 2.5%.
TABLE II: Feature importance scores of the FSR sensors
@@
The above results suggest that some sensors can be removed without a significant compromise in accuracy. To further evaluate this, we progressively remove sensors with importance scores below specific thresholds and evaluate the accuracy of a newly trained model with the retained sensors. In the results presented in Table III, eliminating low-scoring sensors is shown to insignificantly affect accuracy, and moderately high thresholds preserve acceptable performance while reducing the overall sensor count. Hence, a more compact FMG device can be developed, potentially lowering both hardware costs and system complexity.
TABLE III: Sensor reduction impact on accuracy
Importance Sensors Error (mm)
threshold (%) retained Wrist Elbow
0.5 28 121 $\pm$ 66 59 $\pm$ 34
1.0 21 127 $\pm$ 67 64 $\pm$ 36
1.5 14 138 $\pm$ 78 71 $\pm$ 41
2.0 12 140 $\pm$ 77 72 $\pm$ 39
2.5 11 147 $\pm$ 81 76 $\pm$ 42
@@ III-D Model Transfer
To evaluate the model’s generalization capabilities, we tested it on three new participants with different Chest Circumferences (CC), Shoulder Widths (SW) and Upper-Arm Circumferences (UAC), who were not included in the original training dataset. We evaluate the zero-shot (ZS) transfer to the new participant and fine-tuning (FT) with a limited amount of data. For the FT, each new participant contributed 20,000 samples over two sessions and an additional 10,000 labeled samples for testing, with the same procedure outlined in Section II-C. FT for a new participant is performed by only retraining the model’s last layers of the FC networks, with a learning rate of $10^{-5}$. Table V summarizes the results along with anthropometric measures of all participants including the one used for training. While the model’s accuracy decreased slightly when tested on new participants in ZS with different anthropometric measurements, the results demonstrate its ability to generalize and perform effectively. To improve transferability, a limited amount of additional data from the new participant is shown to significantly improve accuracy.
TABLE IV: Accuracy of model transfer to new users
@@
ZS FT ZS FT
Train 98 44 28 60 $\pm$ 60 - 120 $\pm$ 33 -
1 99 46 30.5 100 $\pm$ 57 82 $\pm$ 34 190 $\pm$ 90 139 $\pm$ 63
2 92 44 25.5 93 $\pm$ 41 59 $\pm$ 31 178 $\pm$ 73 120 $\pm$ 62
3 96 50 34.5 99 $\pm$ 42 69 $\pm$ 37 178 $\pm$ 79 122 $\pm$ 65
TABLE V: HRI demonstration success rates
@@ III-E Real-time HRI Demonstration
Two experimental sessions were conducted. In the first session, the user deliberately interfered the motion of the robot by blocking its path. An example of such trial is demonstrated in Figure 7. In the second session, the user naturally performed various tasks, such as picking up and placing tools and other objects scattered on the table (Figures 8-9), while blocking the path of the robot. Each of the sessions consisted of 60 trials. Table V presents the success rate for the two sessions, including metrics for false positives with failure to stop and false negatives with unnecessary stops. Reasons for false positives and negatives may stem from momentary large errors or out-of-distribution motions, such as high accelerations, which caused the robot to incorrectly estimate a clear path or a potential collision, respectively. Nevertheless, the majority of trials resulted in true positives, demonstrating the potential of this approach for natural HRI, without visual perception.
@@ IV Conclusions
We investigated the use of a low-cost wearable FMG device for human arm pose estimation in Human-Robot Interaction (HRI) scenarios. This device, worn on the user’s shoulder, incorporates 32 FSR sensors. Using a Transformer-based model, FMG measurements are mapped to the positions of the elbow and wrist, with respect to the shoulder. Our approach exhibited the ability to provide robust, environment-agnostic and fairly accurate arm pose estimations. Unlike common IMU solutions for arm pose estimation, the FMG device can provide drift-free and absolute estimations without the need for frequent calibrations. The trained model was also shown to be transferable to new users with slight decline in accuracy. Furthermore, a set of HRI experiments has demonstrated the robot’s ability to avoid collisions using the estimated arm poses in a shared workspace. Generally, the FMG device has the potential to enhance various HRI applications, including medical procedures, factory work and assistive technologies.
While the estimation model was shown to be able to fairly generalize to new users, future work should consider making the model more robust to new users. One approach to enhance the model’s generalization capabilities is to train it on a diverse dataset encompassing data from multiple users, as demonstrated in [31]. Alternatively, future work could involve customizing both the FMG device and the model to individual users. This could include 3D scanning a user’s body to create a custom-fitted device, ensuring optimal sensor placement and data quality. Then, the model can be customized for the user through fine-tuning with a limited amount of additional data, as in [26].
To further enhance accuracy, future work could extend the model to estimate limb orientation. This additional information could help mitigate error propagation and refine wrist position estimations. Furthermore, limb orientation estimates could be valuable for tasks such as human-robot hand-over interactions. These suggestions have the potential to also mitigate false positive and negative readings, while future work could also expand the training dataset to include complex and high-velocity motions, thereby enabling the system to handle rapid and less common movements effectively. Unlike EMG, FMG can be placed-on or integrated into the fabric of clothing, enhancing comfort and user experience. Hence, further work can integrate FSR sensors into wearable textiles and open up new possibilities for advanced and more spontaneous HRI. Future work could also explore the integration of FMG with vision to create a robust hybrid system capable of handling diverse environmental conditions. FMG can provide robust signals unaffected by visual uncertainties, while visual perception can enhance spatial awareness and localize general human pose within the environment.